# Theoretical Studies of Sub-THz Active Split-Ring Resonators for Near-Field Imaging

### v1

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Near-fields enable high-resolution sub-wavelength imaging systems by overcoming the fundamental resolution limitations of far-field imaging, which is constrained to approximately half the wavelength of the excitation signal [huang_diffraction_cell2010]. Unlike far-fields, near-fields can be conveniently generated on a chip surface using compact, pixelated designs. The most common approaches include coplanar electrode arrays at RF and microwave frequencies [hu_imgr_tbiocas2022, cheng_imgr_mwtl2023] and resonant structures at mm-wave frequencies and above [mitsunaka_jssc2016, hillger_jssc2018, tanaka_tcas2022].
When the goal is to image biological cells and tissues, at lower frequencies, ionic and interfacial polarization sensing takes place. However, at higher frequencies, these low-frequency barriers are bypassed, providing access to intracellular information and cytoplasmic changes [Abasi_cellsens_acsmeas2022]. Therefore, the extreme sensitivity of mm-wave and THz signals to water content can be leveraged as a natural biomarker for non-invasive imaging of tissue malignancies [Pfeiffer_uwavemag2019, cassar_thzimg_scirep2021]. These signals further benefit from their shorter wavelengths, enabling a smaller pixel size and, thus, high-resolution imaging. Furthermore, the narrowband nature of resonators results in improved noise performance relative to wideband sensing. The combination of these distinctive characteristics has motivated researchers to develop high-resolution, high-contrast resonator-based mm-wave and THz near-field imaging systems.
@@
The architectures proposed in [ameri_vlsi2019, mitsunaka_jssc2016, tanaka_tcas2022], as shown in Fig. 1(1(a)), were limited to using a single detector and a single signal source per imaging pixel. While these configurations can achieve high sensitivity, the pixel size is determined not only by the resonator dimensions but also by the additional components of the source and detector. Therefore, the imaging resolution does not scale consistently with the frequency of operation.
A more effective approach was introduced in [Grzyb_tmtt2017, hillger_jssc2018], where pixels were coupled to, rather than directly connected to, the excitation source and detector. This allowed for physical separation between these components. This approach, as shown in Fig. 1(1(b)), enabled a single source to excite multiple pixels. However, each pixel still required its dedicated local detector, limiting the array scalability to a single dimension.
A key limitation of the previously discussed architecture is the inability to enable and disable imaging resonators selectively. This feature is essential for sharing a single source and a single detector among multiple pixels. By doing so, a two-dimensional array can be constructed that not only scales in both dimensions but also scales well with frequency, since each pixel consists solely of a resonator, whose size decreases with increasing frequency. This concept is depicted in Fig. 1(1(c)) and was recently demonstrated in hardware in [ameri_isscc], where animal tissue samples were imaged with unprecedented contrast at tens-of-microns resolution.
The reader is referred to [ameri_isscc] to learn more about the details of the imager architecture. However, the following provides a high-level overview: a sub-THz signal is generated and split into two paths. The majority of the signal power is used as the LO signal for the detector. The remaining power is coupled to the RF path and is used for imaging. The RF path consists of a transmission line loaded with an array of imaging pixels. The sample shifts the phase characteristics of the pixels. These shifts are measured by sequentially enabling the pixels and performing phase detection by mixing the LO and RF signals. An image frame is constructed once all pixels in the array are measured.
@@ II Split-Ring Resonator (SRR) Analysis
The resonator used in this work is a broadside-coupled SRR equipped with a positive feedback circuit, i.e., $g_{m}$, collectively referred to as an ASRR, which serves as a single pixel in the imaging array. The $g_{m}$ block can be switched on and off to alternate between low-Q, $Q_{OFF}$, and high-Q, $Q_{ON}$, resonance modes, a feature essential to array functionality.
In this section, the effects of $g_{m}$ are set aside, and they will be reintroduced into our analyses in Section III. An SRR consists of two concentric rings with a gap opening on each ring. There are multiple design choices for coupling the rings to each other and the placement of the gap on each ring. In applications where the resonators must be as compact as possible, besides increasing the resonance frequency, these choices make a significant difference. In [rmarques_srr_tant2003], two possible coupling modes of the rings, namely edge coupling and broadside coupling, were studied. While both of these coupling modes can be easily implemented on chip, broadside coupling is preferred as it takes advantage of parallel plate capacitances between the two rings, which can be significantly higher compared to the parallel plate and fringing capacitance of the edge-coupled rings, as shown in Fig. 2(a).
Figure 2: Different SRR coupling configurations: (a) Edge-coupled, (b) Broadside-coupled. (c) HFSS EM simulated resonance frequencies of the resonators as a function $\theta$.
@@
Figure 5: The magnitude and phase of $S_{21}$ while changing $k$ for when (5(a)) $Q$ is kept constant, and (5(b)) $Q$ is calculated from (4). (5(c)) The contours of $S_{11}$ for various $(k,Q_{ON})$ pairs and the boundary where $S_{11}<-10dB$. (5(d)) $S_{11}$, $S_{21}$, and $Q_{ON}$ of EM simulated SRR compared to the calculated values as a function of SRR spacing to the host transmission line, and therefore $k$.
@@ II-A Equivalent Circuit Models
The SRR coupled to a lossless transmission line, with a phase constant of $\beta$ and a characteristic impedance of $Z_{0}$, as shown in Fig. 3(3(a)) can be modeled using a distributed LC model for the transmission line and a lumped LC model for the resonator as illustrated in Fig. 3(3(b)). Here, $L_{TL}$ and $C_{TL}$ represent the distributed inductance and capacitance values of the transmission line section of length $l$. As such, $\beta=\omega\sqrt{L_{TL}C_{TL}}/l$ and $Z_{o}=\sqrt{L_{TL}/C_{TL}}$. The SRR itself is modeled via $L_{SRR}$ and $C_{SRR}$, where the losses are modeled with a series $r_{SRR}=\omega_{0}L_{SRR}/Q$. The capacitive couplings between the SRR and the transmission line are negligible compared to the magnetic couplings, which is denoted by $k$. The impedance, $Z_{1}$, of the combination of SRR coupled to the transmission line, assuming that it is terminated by $Z_{o}$, is given by
@@
where $\omega_{0}=1/{\sqrt{L_{SRR}C_{SRR}}}=1/{\sqrt{{L^{{}^{\prime}}}_{SRR}{C^{{}^{\prime}}}_{SRR}}}$ is the SRR resonance frequency. This equivalent model is depicted in Fig. 3(3(c)). These calculations supplement the results presented in [fmartin_appphylet2003, baena_srr_tmtt2005, shang_srr_tmtt2013] using other methods for lossless resonators. Including the loss has numerous implications in the design for a proper input match, as well as sensitivity and noise analyses. Around $\omega_{0}$, the circuit reduces to Fig. 3(3(d)), which will be used to derive the optimum coupling conditions subsequently.
@@ II-B Coupling Effects and Optimum Coupling
Equation (4) plays a crucial role in the subsequent analyses and has been verified through simulations. In the EM model presented earlier, $k$ is varied by changing the space between the resonator and the transmission line, denoted by the variable $S_{SRR}$, in Fig. 5(5(d)). The value of $k$ for each $S_{SRR}$ was determined through curve-fitting of the LC model response to that of the EM model. $\beta$ can also be derived from the EM model by characterizing the host transmission line, and $l$ is the length of the segment of the line that interacts with the resonator. Then, the requisite $Q_{ON}$ was calculated from (4) and was realized using an ideal negative resistance placed in parallel with the SRR. The calculated and simulated $Q_{ON}$ values to achieve optimum coupling are plotted against each other in Fig. 5(5(d)). More importantly, this figure also contains the resulting return loss and forward transmission for different coupling configurations. It can be seen that an $S_{11}\approx-10dB$ and $S_{21}\approx-4dB$ are consistently achieved as predicted by the theory, which confirms the validity of equation (4), as well as, once again, the utility of the LC model.
@@ II-C Signal Detection
The foregoing discussion indicates that a parallel LC resonance response is expected from an SRR, characterized by a notch in the amplitude and a high-slope region in the phase of the output signal. This provides the opportunity for both power and phase detection. Power detection, due to its wideband operation and compactness, is a widely used technique in high-frequency sensing and spectroscopy systems, such as in [hillger_jssc2018, wang_gasspec_jssc2021, laemmle_sens_tmtt2013]. However, the high phase noise of sub-THz VCOs can severely compromise the detection SNR. Alternatively, phase detection provides immunity to the VCO phase noise. As will be shown in Section IV-D, the pixel input and output close-in phase noise are identical, and thus the phase noise, as a correlated signal, gets suppressed by phase detection. Since each pixel operates in a very narrow band around its resonance frequency, a quadrature phase relationship between the $LO$ and $RF$ signals can be established by introducing an adequate phase shift in one of these signal paths. Consequently, the output of the phase detector can be expressed as
@@ III Active SRR (ASRR) Analysis
The introduction of the sample to an ASRR changes both $C_{SRR}$ and $R_{SRR}$. These changes can be expressed as $\Delta C=f^{{}^{\prime}}(\Delta\Re({\varepsilon}))$ and $\Delta R=f^{{}^{\prime\prime}}(\Delta\Im({\varepsilon}))$, as studied in [Ferrier_loc, Elhadidy_sens_tcas2015, chien_reactsens_jssc2016], where $\varepsilon=\Re({\varepsilon})-j\Im({\varepsilon})$ is the complex permittivity of the sample. In the case of passive SRR, $\Delta C$ and $\Delta R$ directly alter the resonance frequency and quality factor, respectively. However, in an ASRR, although $\Delta C$ still directly changes the resonance frequency, the effect of $\Delta R$ is amplified by the Q-boosting factor squared, using (11) and (12) one arrives at
@@
The circuit in Fig. 7(7(a)) with the component values summarized in the table was used to verify the above derivations. $C_{SRR}$ and $R_{SRR}$ were varied while measuring the resonance frequency and $S_{RES}$. Based on (15), $\Delta\omega_{0}/\Delta C_{SRR}=-5.35\times 10^{25}$ (rad/s.F), which matches the slope of the curve shown in Fig. 7(7(b)).
@@ IV-A Device White Noise
Fig. 12(12(a)) shows the simulation results of output noise density for an optimally coupled ASRR, for $g_{m}$ built from square-law and PDK devices, compared to theoretical values derived from (31), where the $\gamma$ value was extracted from simulating a single device. Since the ASRR is operating in the linear regime and the input signal does not modulate the white noise originating from the ASRR, the noise from (31) is additive and thus splits evenly between AM and PM [Darabi_rfic]. Therefore, the SSB phase noise can be calculated by taking the ratio between the PM sideband power and the carrier power as follows
@@ IV-B Device Flicker Noise
The low-frequency noise of the devices affects the resonator response in an entirely different way. Both $\omega_{0}$ and $Q_{ON}$ can be modulated by changing the DC operating points of the negative resistance due to $1/f$ noise. The former is caused by variations in the device parasitic capacitances, which are quite negligible and will not be considered in our discussions. The latter, however, can be more significant and warrants further study.
@@
which result in a change in the transconductance of the devices, denoted by $dg_{mi}\,(i=1,2,3,4)$. Furthermore, for two transconductances in series after undergoing two small arbitrary changes, $\delta_{1}$ and $\delta_{2}$, the total transconductance can be estimated from
@@
Equation (39) was verified by simulations while sweeping $Q_{ON}$, applying a small offset $v_{fn}$, and measuring the ratio, $dS_{RES}/v_{fn}$. The results are shown in Fig. 13(13(b)), where the theoretical values from (39) are compared to simulation results for square-law and CMOS 28nm PDK devices. It is noteworthy that the output noise exhibits a quadratic dependence on $Q_{ON}$, posing a challenge if excessively large boosted quality factors are targeted.
@@ IV-C Supply Noise
Assuming that the supply is provided by an off-chip ultra-low-noise LDO, such as ADI’s LT3045-1, simulations show that the phase noise caused by the supply is negligible compared to the noise generated by the $g_{m}$ devices and thus does not merit further attention.
@@ IV-D Input Signal Phase Noise
It is evident that the noise on a signal propagating through a transmission line, in the absence of the resonator, directly appears at the output, albeit with some delay. However, as illustrated in Fig. 14(14(a)), in the presence of a resonator, the output phase accumulates an extra phase, which depends on the rate at which the input phase changes. In other words, a sudden shift, $\varphi_{n}$, due to the input phase noise not only shows up in the output phase, but also gives rise to an extra phase shift through the resonator, resulting in a total output phase error of
@@
Figure 14: (14(a)) Phase noise propagation through an unloaded transmission line and a loaded transmission line with a resonator. (14(b)) Simulations of the input and output phase noise of a loaded transmission line.
@@
A signal traveling through a resonator experiences PM-to-AM conversion, which can also be derived analytically. This conversion peaks at the inflection points of the magnitude of $S_{21}$ and goes to zero at the resonance frequency. As pointed out in [wang_gasspec_tbiocas2018], the two PM sidebands experience an imbalance after interacting with the resonator, as shown in Fig. 15(15(a)). The resulting sidebands can be decomposed into two pairs of PM and AM sidebands, where the PM sidebands are roughly identical to those of the input, as discussed above, and the AM sidebands are thus given by half the difference between the output sidebands. Assuming the input signal contains purely PM sidebands, utilizing narrow-band approximation, they can be expressed as
@@
where the second and third terms on the right-hand side are lower and upper sidebands (LSB and USB), respectively, and $\theta$ is an arbitrary phase difference between the fundamental tone and the sidebands. The average of the phases of the two sidebands is $\pi/2$, making their summation perpendicular to the fundamental tone, thus only modulating the phase [Darabi_rfic]. Additionally, $\Delta\varphi A/2$ captures the phase noise at a frequency offset $\Delta\omega$, whose power normalized to the fundamental tone power is the input phase noise, $\mathcal{L}_{\varphi}(\Delta\omega)$.

### v2

@@ Abstract
Index Terms:
Active split-ring resonator, sub-THz, near-field, quality factor boosting, sensing, imaging.
© 2026 IEEE. All rights reserved, including rights for text and data mining and training of artificial intelligence and similar technologies. Personal use is permitted, but republication/redistribution requires IEEE permission. See https://www.ieee.org/publications/rights/index.html for more information
@@ I Introduction
Near-fields enable high-resolution sub-wavelength imaging systems by overcoming the fundamental resolution limitations of far-field imaging, which is constrained to approximately half the wavelength of the excitation signal [14]. Unlike far-fields, near-fields can be conveniently generated on a chip surface using compact, pixelated designs. The most common approaches include coplanar electrode arrays at RF and microwave frequencies [13, 6] and resonant structures at mm-wave frequencies and above [19, 12, 22].
When the goal is to image biological cells and tissues, at lower frequencies, ionic and interfacial polarization sensing takes place. However, at higher frequencies, these low-frequency barriers are bypassed, providing access to intracellular information and cytoplasmic changes [1]. Therefore, the extreme sensitivity of mm-wave and THz signals to water content can be leveraged as a natural biomarker for non-invasive imaging of tissue malignancies [20]. These signals further benefit from their shorter wavelengths, enabling a smaller pixel size and, thus, high-resolution imaging. Furthermore, the narrowband nature of resonators results in improved noise performance relative to wideband sensing. The combination of these distinctive characteristics has motivated researchers to develop high-resolution, high-contrast resonator-based mm-wave and THz near-field imaging systems.
@@
The architectures proposed in [4, 19, 22], as shown in Fig. 1(1(a)), were limited to using a single detector and a single signal source per imaging pixel. While these configurations achieve high sensitivity, the pixel size is determined not only by the resonator dimensions but also by the additional components of the source and detector. Therefore, the imaging resolution does not scale consistently with the frequency of operation.
A more effective approach was introduced in [11, 12], where pixels were coupled to, rather than directly connected to, the excitation source and detector. This allowed for separation among these components. This approach, as shown in Fig. 1(1(b)), enabled a single source to excite multiple pixels. However, each pixel still required its dedicated local detector, limiting the array scalability to a single dimension.
A key limitation of the discussed architecture is the inability to enable and disable imaging resonators selectively. This feature is essential for sharing a single source and a single detector among multiple pixels. By doing so, a two-dimensional array can be constructed that not only scales in both dimensions but also scales well with frequency, since each pixel consists solely of a resonator, whose size decreases with increasing frequency. This concept is depicted in Fig. 1(1(c)) and was demonstrated in [3], where animal tissue samples were imaged with high contrast at tens-of-microns resolution.
The reader is referred to [2] for a detailed discussion of the imager architecture and imaging experiments. However, the following provides a brief overview: a sub-THz signal is generated and split into two paths. The majority of the signal power is used as the LO signal for the detector. The remaining power is coupled to the RF path and is used for imaging. The RF path consists of a transmission line loaded with an array of imaging pixels. The sample shifts the phase characteristics of the pixels. These shifts are measured by sequentially enabling the pixels and performing phase detection by mixing the LO and RF signals. An image frame is constructed once all pixels are measured.
@@ II Split-Ring Resonator (SRR) Analysis
The resonator used in this work is a broadside-coupled SRR equipped with a positive feedback circuit, i.e., $g_{m}$, collectively referred to as an ASRR, which serves as a single pixel in the imaging array. As shown in Fig. 6(6(b)), the $g_{m}$ block can be switched on and off to alternate between low-$Q$, $Q_{OFF}$, and high-$Q$, $Q_{ON}$, resonance modes, a feature essential to array functionality.
In this section, the effects of $g_{m}$ are set aside, and they will be reintroduced into our analyses in Section III. An SRR consists of two concentric rings with a gap opening on each ring. For understanding SRR EM behavior, the reader is referred to [17]. There are multiple design choices for coupling the rings to each other and the placement of the gap on each ring. In imaging applications where the resonators must be as compact as possible, these layout choices, besides increasing the resonance frequency, can significantly reduce the resonator dimensions and thus increase the imaging resolution. In [16], two possible coupling modes of the rings, namely edge coupling and broadside coupling, were studied. While both of these coupling modes can be easily implemented on chip, broadside coupling is preferred as it takes advantage of parallel plate capacitances between the two rings, which can be significantly higher compared to the parallel plate and fringing capacitance of the edge-coupled rings, as shown in Fig. 2(a).
Figure 2: (a) Edge-coupled SRR, (b) Broadside-coupled SRR. (c) HFSS simulated resonance frequency as a function of $\theta$.
@@
Figure 5: The magnitude and phase of $S_{21}$ while sweeping $k$ for (5(a)) constant $Q$, and (5(b)) calculated $Q$ from (4), where $\beta l=0.25$. (5(c)) The contours of $S_{11}$ for various $(k,Q_{ON})$ pairs and the boundary where $S_{11}<-10dB$. (5(d)) $S_{11}$, $S_{21}$, and $Q_{ON}$ of EM simulated SRR compared to the calculated values as a function of SRR spacing to the host transmission line, and therefore $k$.
@@ II-A Equivalent Circuit Models
The SRR coupled to a lossless transmission line, with a phase constant of $\beta$ and a characteristic impedance of $Z_{0}$, as shown in Fig. 3(3(a)) can be modeled using a distributed LC model for the transmission line and a lumped LC model for the resonator as illustrated in Fig. 3(3(b)). Here, $L_{TL}$ and $C_{TL}$ represent the distributed inductance and capacitance values of the transmission line section of length $l$. As such, $\beta=\omega\sqrt{L_{TL}C_{TL}}/l$ and $Z_{o}=\sqrt{L_{TL}/C_{TL}}$. The losses of the transmission line can be neglected as $l$ is small (on the order of the SRR length, which is $\sim$30-40 $\mu m$). The SRR is modeled via $L_{SRR}$ and $C_{SRR}$, where the losses are modeled with $r_{SRR}=\omega_{0}L_{SRR}/Q$. The capacitive couplings between the SRR and the transmission line are negligible compared to the magnetic couplings, which is denoted by $k$. The impedance, $Z_{1}$, of the combination of SRR coupled to the transmission line, assuming that it is terminated by $Z_{o}$, is given by
@@
where $\omega_{0}=1/{\sqrt{L_{SRR}C_{SRR}}}=1/{\sqrt{{L^{{}^{\prime}}}_{SRR}{C^{{}^{\prime}}}_{SRR}}}$ is the SRR resonance frequency. This equivalent model is depicted in Fig. 3(3(c)). These calculations supplement the results presented in [18, 5, 21] using other methods for lossless resonators. Including the loss has numerous implications in the design for a proper input match, as well as sensitivity and noise analyses. Around $\omega_{0}$, the circuit reduces to Fig. 3(3(d)), which will be used to derive the optimum coupling conditions subsequently.
@@ II-B Coupling Effects and Optimum Coupling
Equation (4) plays a crucial role in the subsequent analyses and has been verified through simulations. In the EM model presented earlier, $k$ is varied by changing the space between the resonator and the transmission line, denoted by the variable $S_{SRR}$, in Fig. 5(5(d)). The value of $k$ for each $S_{SRR}$ was determined through curve-fitting of the LC model response to that of the EM model. $\beta$ can also be derived from the EM model by characterizing the host transmission line, and $l$ is the length of the segment of the line that interacts with the resonator. Then, the requisite $Q_{ON}$ was calculated from (4) and was realized using an ideal negative resistance placed in parallel with the SRR. The calculated and simulated $Q_{ON}$ values to achieve optimum coupling are plotted against each other in Fig. 5(5(d)). More importantly, this figure also contains the resulting return loss and forward transmission for different coupling configurations. It can be seen that an $S_{11}<-9.25dB$ and $S_{21}>-4dB$ are consistently achieved as predicted by the theory, which confirms the validity of (4) and the utility of the LC model.
@@ II-C Signal Detection
The foregoing discussion indicates that a parallel LC resonance response is expected from an SRR, characterized by a notch in the amplitude and a high-slope region in the phase of the output signal. This provides the opportunity for both power and phase detection. Power detection, due to its wideband operation and compactness, is a widely used technique in high-frequency sensing and spectroscopy systems, such as in [12, 24, 15]. However, the high phase noise of sub-THz VCOs can severely compromise the detection SNR. Alternatively, phase detection provides immunity to the VCO phase noise. As will be shown in Section IV-D, the pixel input and output close-in phase noise are identical, and thus the phase noise, as a correlated signal, gets suppressed by phase detection. Since each pixel operates in a very narrow band around its resonance frequency, a quadrature phase relationship between the $LO$ and $RF$ signals can be established by introducing an adequate phase shift in one of these signal paths. Consequently, the output of the phase detector can be expressed as
@@ III Active SRR (ASRR) Analysis
The introduction of the sample to an ASRR changes both $C_{SRR}$ and $R_{SRR}$. These changes can be expressed as $\Delta C=f^{{}^{\prime}}(\Delta\Re({\varepsilon}))$ and $\Delta R=f^{{}^{\prime\prime}}(\Delta\Im({\varepsilon}))$, as studied in [10, 9, 7], where $\varepsilon=\Re({\varepsilon})+j\Im({\varepsilon})$ is the complex permittivity of the sample. In the case of passive SRR, $\Delta C$ and $\Delta R$ directly alter the resonance frequency and quality factor, respectively. However, in an ASRR, although $\Delta C$ still directly changes the resonance frequency, the effect of $\Delta R$ is amplified by the Q-boosting factor squared, using (11) and (12) one arrives at
@@
The circuit in Fig. 7(7(a)) with the component values summarized in the table was used to verify these derivations. $C_{SRR}$ and $R_{SRR}$ were varied while measuring $\omega_{0}$ and $S_{RES}$. Based on (15), $\Delta\omega_{0}/\Delta C_{SRR}=-5.35\times 10^{25}$ (rad/s.F), which matches the slope of the curve shown in Fig. 7(7(b)).
@@ IV-A Device White Noise
Fig. 12(12(a)) shows the simulation results of output noise density for an optimally coupled ASRR, for $g_{m}$ built from square-law and PDK devices, compared to theoretical values derived from (31), where the $\gamma$ value was extracted from simulating a single device. Since the ASRR is operating in the linear regime and the input signal does not modulate the white noise originating from the ASRR, the noise from (31) is additive and thus splits evenly between AM and PM [8]. Therefore, the SSB phase noise can be calculated by taking the ratio between the PM sideband power and the carrier power as follows
@@ IV-B Device Flicker Noise
The low-frequency noise of the devices affects the resonator response differently. Both $\omega_{0}$ and $Q_{ON}$ can be modulated by changing the DC operating points of the negative resistance due to $1/f$ noise. The former is caused by variations in the device parasitic capacitances, which are quite negligible and will not be considered in our discussions. The latter, however, can be more significant and warrants further study.
@@
which result in a change in device transconductance, denoted by $dg_{mi}\,(i=1,2,3,4)$. Furthermore, for two transconductances in series after undergoing two small arbitrary changes, $\delta_{1}$ and $\delta_{2}$, the total transconductance can be estimated from
@@
Equation (39) was verified by simulations while sweeping $Q_{ON}$, applying a small offset $v_{fn}$, and measuring the ratio, $dS_{RES}/v_{fn}$. The results are shown in Fig. 13(13(b)), where the theoretical values from (39) are compared to simulation results for square-law and CMOS 28nm PDK devices. The output noise exhibits a quadratic dependence on $Q_{ON}$, posing a challenge if large boosted quality factors are targeted.
@@ IV-C Supply Noise
If the supply is provided by an off-chip low-noise LDO, such as ADI’s LT3045-1, simulations show that the phase noise caused by the supply is negligible compared to device noise contribution and thus does not merit further attention.
@@ IV-D Input Signal Phase Noise
The noise on a signal propagating through a transmission line, in the absence of the resonator, directly appears at the output, albeit with some delay. However, as illustrated in Fig. 14(14(a)), in the presence of a resonator, the output phase accumulates an extra phase, which depends on the rate at which the input phase changes. In other words, a sudden shift, $\varphi_{n}$, due to the input phase noise not only shows up in the output phase, but also gives rise to an extra phase shift through the resonator, resulting in a total output phase error of
@@
Figure 14: (14(a)) Phase noise propagation through a transmission line and an SRR-loaded transmission line. (14(b)) Simulations of the input and output phase noise of a loaded transmission line.
@@
A signal traveling through a resonator experiences PM-to-AM conversion, which can also be derived analytically. This conversion peaks at the inflection points of the magnitude of $S_{21}$ and goes to zero at the resonance frequency. As pointed out in [23], the two PM sidebands experience an imbalance after interacting with the resonator, as shown in Fig. 15(15(a)). The resulting sidebands can be decomposed into two pairs of PM and AM sidebands, where the PM sidebands are roughly identical to those of the input, as discussed above, and the AM sidebands are thus given by half the difference between the output sidebands. Assuming the input signal contains purely PM sidebands, utilizing narrow-band approximation, they can be expressed as
@@
where the second and third terms on the right-hand side are lower and upper sidebands (LSB and USB), respectively, and $\theta$ is an arbitrary phase difference between the fundamental tone and the sidebands. The average of the phases of the two sidebands is $\pi/2$, making their summation perpendicular to the fundamental tone, thus only modulating the phase [8]. Additionally, $\Delta\varphi A/2$ captures the phase noise at a frequency offset $\Delta\omega$, whose power normalized to the fundamental tone power is the input phase noise, $\mathcal{L}_{\varphi}(\Delta\omega)$.